# Privacy-Preserving Instructions for Aligning Large Language Models

### v1

@@ D.3 Run Algorithm 2 with Fewer Initial Synthetic Instructions
D.4 Evaluate SFT Models on AlpacaEval
@@
In addition to the test set of Chatbot Arena instructions, we also evaluate the models in Section 4.3 on the AlpacaEval dataset (Dubois et al., 2023). AlpacaEval comprises 805 instructions collected from public evaluation sets, including MT-Bench (Zheng et al., 2023c), the self-instruct evaluation set (Wang et al., 2022), the Koala evaluation set (Geng et al., 2023), etc. To prevent data contamination, we again apply an 8-gram deduplication between AlpacaEval instructions and Chatbot Arena instructions. We note that AlpacaEval instructions do not accurately represent real-world user queries to LLMs. For example, the self-instruct evaluation set is generated via prompting GPT-3. Nonetheless, we include the results for a more comprehensive evaluation. We use the same evaluation setup as that used by the AlpacaEval leaderboard [^3] https://tatsu-lab.github.io/alpaca_eval/, which allows us to make direct comparisons with state-of-the-art models listed on the leaderboard. The results are presented in Table 6. When evaluated on AlpacaEval instructions, the best performing DP models slightly underperform Vicuna-v1.3 but are comparable with LLaMA2-Chat.
@@ Appendix E Data Cleansing for the LMSYS-Chat-1M Dataset
This section details the data pre-processing procedures applied to the LMSYS-Chat-1M dataset (Zheng et al., 2023a). It is worth noting that for the analysis of sensitive information in user instructions, we utilize a separate dataset: the Chatbot Arena Conversations dataset [^4] https://huggingface.co/datasets/lmsys/chatbot_arena_conversations. The pre-processing steps outlined below are not applied to it.
@@
Take single-round instructions from English conversations. We first extract user instructions from the dataset, while discarding the responses from LLMs. For the purpose of simplifying our experimental setup, we only consider the instruction from the first round in case of multi-round conversations. Zheng et al. (2023a) provide a language tag for each conversation. We only take instructions from English conversations as English accounts for around 80% of the instructions among over more than twenty languages. Approximately 780 thousand instructions remain after this step.
Remove inappropriate instructions. For each conversation in LMSYS-Chat-1M, Zheng et al. (2023a) utilize OpenAI’s moderation API to provide a tag that indicates if a conversation contains inappropriate content. Additionally, they redact real person names in the conversations. We exclude conversations that are either flagged by the moderation API or contain redacted names. Approximately 497 thousand instructions remain after this step.
Sequence de-duplication. Users might repeatedly ask the same question to elicit responses from different LLMs. If an instructions appears multiple times, we only keep it once. There are around 310 thousand instructions left after this step.
N-gram de-duplication. We divide the instructions into training, validation, and test sets using random splits. To prevent data contamination, we implement an 8-gram de-duplication process, aligning with the threshold established by Radford et al. (2019). During this process, if an 8-gram is present in multiple instructions, then only the first instruction that has it will be kept. As a result, the dataset is reduced to approximately 232 thousand instructions. Additionally, since we use the Alpaca Eval instructions (Dubois et al., 2023) as one of our evaluation sets, we further exclude any instructions containing 8-grams that appear in the Alpaca Eval instructions. This leads to the removal of fewer than twenty instructions.
@@
After removing those unusual repetitions, there are around 210K instructions left. We then randomly divide these remaining instructions into three subsets: 180 thousand for training, 10 thousand for validation, and 20 thousand for testing.
@@ F.1 Setup for Computing MAUVE
We use the official implementation [^5] https://github.com/krishnap25/mauve to compute MAUVE scores. For computing MAUVE with Sentence-T5 embeddings, we follow the default setup to quantize the embeddings into 500 bins. The absolute MAUVE scores are affected by a scaling constant $c$. Using a larger $c$ would result in a lower score. For computing MAUVE with unigram distributions, we use the default setup $c=5$. For computing MAUVE with Sentence-T5 embeddings, we set $c=10$ to better distinguish the scores of different datasets. We note that values of $c$ do not affect the ranking of different datasets.
@@ F.3 Hyperparameters for Sampling
For supervised fine-tuning, we follow the prompt for Vicuna. The prompt template is ‘A chat between a curious user and an artificial intelligence assistant. The assistant gives helpful, detailed, and polite answers to the user’s questions. USER: [Instruction] ASSISTANT:’. The maximum sequence length for the generated responses is set as 2048 tokens. During inference, we follow the default setup of Vicuna [^6] https://github.com/lm-sys/FastChat/tree/main. Specifically, sampling temperature $T$ and probability cutoff parameter top_pp are set as 0.7 and 1.0, respectively.
@@ F.4 Code Snippet for Privacy Accounting
We use gpt-3.5-turbo-0301777https://platform.openai.com/docs/models/gpt-3-5 to create responses for all instructions. We use 20 parallel threads to query the OpenAI API. Annotating 100 thousand instructions takes approximately 12 hours and \$80. The maximum response length is 2048 tokens. The sampling temperature $T$ is set as 0.7. The probability cutoff parameter top_pp is set to 0.95.
@@ F.6 Hyperparameters for Proximal Policy Optimization
The baseline model [^8] https://huggingface.co/Open-Orca/oo-phi-1_5 is a 1.3B model fine-tuned on the OpenOrca dataset. The reward model [^9] https://huggingface.co/OpenAssistant/reward-model-deberta-v3-large-v2 comprises 304M parameters and is trained on several public human preference datasets, such as the annotated version of the TL;DR dataset featuring Reddit posts (Völske et al., 2017; Stiennon et al., 2020). The maximum sequence length for generated responses is set as 256 tokens. The sampling temperature and the probability cutoff parameter top_pp are both set to 1.0 during training. Our PPO implementation is based on the Transformer Reinforcement Learning (TRL) library. We follow the default setup for PPO training [^10] https://github.com/huggingface/trl/blob/main/trl/trainer/ppo_config.py for most of the hyperparameters. The KL penalty coefficient is set as 0.2 and the number of optimisation steps per batch is 4. We use a batchsize of 128 and a learning rate of $1.41\times 10^{-5}$. All models are trained for 2500 batches of samples.
@@ Appendix G Additional Figures
We put some figures here because of the space constraints. In Figure 16, we show one demonstration in the prompt for GPT-3.5-Turbo. In Figure 17, we present random samples from two clusters of initial synthetic samples: one cluster receiving the highest number of votes from real data and the other the lowest. In Figure 18, we compare the sequence length distribution of instructions with that of responses from GPT-3.5-Turbo.
INSTRUCTION = "Subject: California Gas Demand Growth
Body: ----------------------
Forwarded by Phillip K Allen/HOU/ECT on 02/12/2001 11:09 AM ---------------------------
From: Mark Whitt@ENRON on 02/09/2001 03:38 PM MST Sent by: Mark Whitt@ENRON
To: Barry Tycholiz/NA/Enron@ENRON, Phillip K Allen/HOU/ECT@ECT, Mike Grigsby/HOU/ECT@ECT, Paul T Lucci/NA/Enron@Enron, Kim Ward/HOU/ECT@ECT, Stephanie Miller/Corp/Enron@ENRON
$c$:
Subject: California Gas Demand Growth
This should probably be researched further.
If they can really build this many plants it could have a huge impact on the Cal border basis. Obviously it is dependent on what capacity is added on Kern, PGT and TW but it is hard to envision enough subscriptions to meet this demand. Even if it is subscribed it will take 18 months to 2 years to build new pipe therefore the El Paso 1.2 Bcf/d could be even more valuable.
Mark
----- Forwarded by Mark Whitt/NA/Enron on 02/09/2001 03:04 PM -----
Tyrell Harrison@ECT
Sent by: Tyrell Harrison@ECT 02/09/2001 09:26 AM
To: Barry Tycholiz/NA/Enron@ENRON, Mark Whitt/NA/Enron@Enron, Paul T Lucci/NA/Enron@Enron, Kim Ward/HOU/ECT@ECT, Stephanie Miller/Corp/Enron@ENRON, Phillip K Allen/HOU/ECT@ECT
$c$:
Subject: California Gas Demand Growth
If you wish to run sensitivities to heat rate and daily dispatch, I have attached the spreadsheet below.
Tyrell 303 575 6478"
RESULT:
[[a]]: [Phillip K Allen, Mark Whitt, Barry Tycholiz, Mike Grigsby, Paul T Lucci, Kim Ward, Stephanie Miller, Tyrell Harrison]
[[b]]: [303 575 6478]
[[c]]: [California, Kern, PGT, TW]
[[d]]: [ENRON]
[[e]]: []
[[f]]: [02/12/2001, 02/09/2001]
Figure 16: One demonstration for initial screening with GPT-3.5-Turbo. The sample is from the Enron Email dataset.

### v2

@@ D.3 Run Algorithm 2 with Fewer Initial Synthetic Instructions
D.4 Run Algorithm 2 on More Datasets
We use two more datasets to show the wide applicability of Algorithm 2. The datasets are 75,000 IMDB movie reviews [^5] https://huggingface.co/datasets/stanfordnlp/imdb. and 75,316 PubMed paper abstracts [^6] https://huggingface.co/datasets/pubmed. (from August 01 to August 05, 2023). To reduce computational cost, the generators are fine-tuned from GPT2-Large. The hyperparameters for DP fine-tuning follow the best ones in Yue et al. (2022). When Algorithm 2 is applied, we increase the noise multiplier for DP fine-tuning to compensate for the privacy cost of the re-sampling process. We generate 750,000 initial synthetic samples and select a subset that is of the same size as real data. Algorithm 2 consistently improves the quality of initial synthetic data, as indicated by the MAUVE scores in Table 6.
Table 6: MAUVE scores (Sentence-T5) of synthetic IMDB reviews and PubMed abstracts. Generators are fine-tuned from GPT2-Large.
IMDB PubMed Synthetic ($\varepsilon=3.96$, no filitering) $0.647$ $0.594$ Synthetic ($\varepsilon=3.93$) $0.825$ $0.811$
D.5 Evaluate SFT Models Using the Non-private Baseline
We evaluate the win rates using a baseline model trained on non-private Chatbot Arena data. The results are presented in Table 7. These findings align with those presented in Table 2 of Section 4.3, where the baseline model is trained on FLAN data. The results show that: 1) the proposed filtering algorithm improves the win-rates; 2) the best-performing DP model are comparable to the non-private baseline.
Table 7: Win-rates with the baseline model trained on non-private Chatbot Arena data.
7B Models Chatbot Arena (non-private) $50\%$ Vicuna-v1.3 $46.0\%$ ($\pm$0.20) Synthetic ($\varepsilon=5.94$, no filt.) $44.2\%$ ($\pm$0.27) Synthetic ($\varepsilon=5.98$) $48.0\%$ ($\pm$0.62) Synthetic (300K, $\varepsilon=5.98$) $49.1\%$ ($\pm$0.51) 13B Models Vicuna-v1.3 $57.7\%$ ($\pm$0.66) Synthetic (300K, $\varepsilon=5.98$) $59.2\%$ ($\pm$0.41)
D.6 Evaluate SFT Models on AlpacaEval
@@
In addition to the test set of Chatbot Arena instructions, we also evaluate the models in Section 4.3 on the AlpacaEval dataset (Dubois et al., 2023). AlpacaEval comprises 805 instructions collected from public evaluation sets, including MT-Bench (Zheng et al., 2023c), the self-instruct evaluation set (Wang et al., 2022), the Koala evaluation set (Geng et al., 2023), etc. To prevent data contamination, we again apply an 8-gram deduplication between AlpacaEval instructions and Chatbot Arena instructions. We note that AlpacaEval instructions do not accurately represent real-world user queries to LLMs. For example, the self-instruct evaluation set is generated via prompting GPT-3. Nonetheless, we include the results for a more comprehensive evaluation. We use the same evaluation setup as that used by the AlpacaEval leaderboard [^7] https://tatsu-lab.github.io/alpaca_eval., which allows us to make direct comparisons with state-of-the-art models listed on the leaderboard. The results are presented in Table 8. When evaluated on AlpacaEval instructions, the best performing DP models slightly underperform Vicuna-v1.3 but are comparable with LLaMA2-Chat.
@@ Appendix E Data Cleansing for the LMSYS-Chat-1M Dataset
This section details the data pre-processing procedures applied to the LMSYS-Chat-1M dataset (Zheng et al., 2023a). For the analysis in Appendix B, we utilize a separate dataset: the Chatbot Arena Conversations dataset [^8] https://huggingface.co/datasets/lmsys/chatbot_arena_conversations. (Zheng et al., 2023c). The pre-processing steps outlined below are not applied to it.
@@
Take single-round instructions from English conversations. We first extract user instructions from the dataset, while discarding the responses from LLMs. For the purpose of simplifying our experimental setup, we only consider the instruction from the first round in case of multi-round conversations. Zheng et al. (2023a) provide a language tag for each conversation. We only take instructions from English conversations as English accounts for around 80% of the instructions among over more than twenty languages. Approximately 780,000 instructions reamain after this step.
Remove inappropriate instructions. For each conversation in LMSYS-Chat-1M, Zheng et al. (2023a) utilize OpenAI’s moderation API to provide a tag that indicates if a conversation contains inappropriate content. Additionally, they redact real person names in the conversations. We exclude conversations that are either flagged by the moderation API or contain redacted names. Approximately 500,000 instructions remain after this step.
Sequence de-duplication. Users might repeatedly ask the same question to elicit responses from different LLMs. If an instructions appears multiple times, we only keep it once. There are around 310,000 instructions left after this step.
N-gram de-duplication. We divide the instructions into training, validation, and test sets using random splits. To prevent data contamination, we implement an 10-gram de-duplication process. During this process, if an 10-gram is present in multiple instructions, then only the first instruction that has it will be kept. As a result, the dataset is reduced to approximately 232 thousand instructions. Additionally, since we use the Alpaca Eval instructions (Dubois et al., 2023) as one of our evaluation sets, we further exclude any instructions containing 10-grams that appear in the Alpaca Eval instructions. This leads to the removal of fewer than twenty instructions. The dataset is reduced to approximately 220,000 instructions after this step.
@@
After removing those unusual repetitions, there are around 200,000 instructions left. We take 180,000 as the training set, 5,000 as the validation set, and the rest as the test set.
@@ Appendix F Implementation Details
This section documents the implementation details of our experiments.
@@ F.1 Setup for Computing MAUVE
We use the official implementation [^9] https://github.com/krishnap25/mauve. to compute MAUVE scores. For computing MAUVE with Sentence-T5 embeddings, we follow the default setup to quantize the embeddings into 500 bins. The absolute MAUVE scores are affected by a scaling constant $c$. Using a larger $c$ would result in a lower score. For computing MAUVE with unigram distributions, we use the default setup $c=5$. For computing MAUVE with Sentence-T5 embeddings, we set $c=10$ to better distinguish the scores of different datasets. We note that values of $c$ do not affect the ranking of different datasets.
@@ F.3 Hyperparameters for Sampling
For supervised fine-tuning, we follow the prompt for Vicuna. The prompt template is ‘A chat between a curious user and an artificial intelligence assistant. The assistant gives helpful, detailed, and polite answers to the user’s questions. USER: [Instruction] ASSISTANT:’. The maximum sequence length for the generated responses is set as 2048 tokens. During inference, we follow the default setup of Vicuna [^10] https://github.com/lm-sys/FastChat/tree/main.. Specifically, sampling temperature $T$ and probability cutoff parameter top_pp are set as 0.7 and 1.0, respectively.
@@ F.4 Code Snippet for Privacy Accounting
We use gpt-3.5-turbo-0301111111https://platform.openai.com/docs/models/gpt-3-5. to create responses for all instructions. We use 20 parallel threads to query the OpenAI API. Annotating 100 thousand instructions takes approximately 12 hours and \$80. The maximum response length is 2048 tokens. The sampling temperature $T$ is set as 0.7. The probability cutoff parameter top_pp is set to 0.95.
@@ F.6 Hyperparameters for Proximal Policy Optimization
The baseline model [^12] https://huggingface.co/Open-Orca/oo-phi-1_5. is a 1.3B model fine-tuned on the OpenOrca dataset. The reward model [^13] https://huggingface.co/OpenAssistant/reward-model-deberta-v3-large-v2. comprises 304M parameters and is trained on several public human preference datasets, such as the annotated version of the TL;DR dataset featuring Reddit posts (Völske et al., 2017; Stiennon et al., 2020). The maximum sequence length for generated responses is set as 256 tokens. The sampling temperature and the probability cutoff parameter top_pp are both set to 1.0 during training. Our PPO implementation is based on the Transformer Reinforcement Learning (TRL) library. We follow the default setup for PPO training [^14] https://github.com/huggingface/trl/blob/main/trl/trainer/ppo_config.py. for most of the hyperparameters. The KL penalty coefficient is set as 0.2 and the number of optimisation steps per batch is 4. We use a batchsize of 128 and a learning rate of $1.41\times 10^{-5}$. All models are trained for 2500 batches of samples.
@@ Appendix G Additional Figures
We put some figures here because of the space constraints. In Figure 17, we present random samples from two particular clusters of initial synthetic instructions: one cluster receiving the highest number of votes from real data and the other receiving the lowest. In Figure 18, we compare the sequence length distribution of Chatbot Arena instructions with the length distribution of responses from GPT-3.5-Turbo.